OpenStudy (is3535):

A $7000 investment earns $1330 simple interest in two years. What is the annual interest rate? A. 1.90% B. 3.80% C. 9.50% D. 19.0%

OpenStudy (is3535):

@shifuyanli

OpenStudy (anonymous):

idk ask @jagr2713

OpenStudy (is3535):

@jagr2713

jagr2713 (jagr2713):

You still here @is3535

OpenStudy (is3535):

yea

jagr2713 (jagr2713):

You'll use the formula I = P*r*t where the interest earned is I = 1330, the amount invested is P = 7000 and the time is t = 2 years. So, I = P*r*t 1330 = 7000*r*2 then you solve for r what do you get @is3535

OpenStudy (is3535):

i think it c

OpenStudy (is3535):

I = Prt This is the formula to find Interest Rate. We are given the Interest Rate, Principal, and time(in years). So we plug in what we know and solve for 'r'.

OpenStudy (is3535):

@sleepyjess

OpenStudy (is3535):

@winsteria

jagr2713 (jagr2713):

CoRRECT